What is the value of 3/4+3/28+3/70+3/130+............+3/9700=?

Grade:11

2 Answers

manjula
8 years ago
 in algebric expression the value of the 3/4+3/28+3/70+3/130+......................................+3/9700 has the answer
 0.923386201428
Sahil Kabra
6 years ago
That answer is incorrect.the correct andwer is 0.99. Here is how:
3/4=1-1/4
3/28=1/4-1/7
3/70=1/7-1/0
…..
3/9700 = 1/97-1/100
Add all the above..What remains is 1-1/100 = .99